EXPERIMENTAL DETERMINATION OF INFRARED TRANSITION DIPOLE MOMENTS FOR HNC FROM HERMAN-WALLIS EFFECT

\underline{M. NEZU} and T. AMANO, Institute for Astrophysics and Planetary Sciences, Ibaraki University, Mito 310, Japan; K. KAWAGUCHI, Nobeyama Radio Observatory, Nobeyama, Nagano 384-13, Japan.

The transition dipole moments for the fundamental bands of HNC were obtained from analyses of Herman-Wallis effect on the absorption intensities. All the fundamental bands were measured using a Fourier transform infrared spectrometer(Bruker IFS 120HR) at Nobeyama Radio Observatory. A glow discharge in a mixture of CH_{3}CN(∼ 50 mTorr), H_{2}(∼ 150 mTorr), and Ar(∼ 100 mTorr) was used for production of HNC. The spectra were recorded with resolution of 0.01 cm^{−1} and the absorption path length was 24 m. The experimentally determined first order Herman-Wallis coefficients for the ν_{1} and ν_{3} bands, combined with the relative value of the transition dipole moments for the ν_{1} and ν_{3} bands derived from the relative intensity measurements, yielded the following values for the transition dipole moments (in Debye), using the expressions for Herman-Wallis coefficients given by Watson^{a},

R_{1} = 0.194(13), R_{2} = −0.886(13), R_{3} = −0.169(11).

These values are considerably larger than the corresponding values for HCN^{b} and are in reasonable agreement with ab initio values^{c}^{d}.

aJ. K. G. Watson, J. Mol.Spectrosc. 125, 428(1987)

bA. G. Maki et al, J. Mol.Spectrosc. 174, 365(1995)

cP. Botschwina et al, Chem. Phys. 190, 345(1995)

dT. J. Lee and A. P. Rendell, Chem. Phys. Lett. 177, 491(1991)